1. The garden strawberry is a widely grown hybrid species of the genus Fragaria, collectively known as the strawberries. It is cultivated worldwide for its fruit. The fruit is widely appreciated for its characteristic aroma, bright red color, juicy texture, and sweetness. It is consumed in large quantiies, either fresh or in prepared foods such as preserves, juice, jam, pies, ice creams, milkshakes, and chocolates. To make strawberry jam, strawberries containing 15 wt% solids and 85 wt% water are crushed. The crushed strawberries and sugar are mixed in a 4/5 mass ratio and the mixture is heated to evaporate water. The residue contains one-third water by mass. Calculate the amounts of strawberries needed to make 100 g of jam, and that of evaporated water. 15 2. In your own words, explain the functions of the following unit operations in chemical engineering processes: (i) quencher, i) decanter, ) turbine, (iv) humidifier, (v) partial condenser 5]
Answer #1

1. Basis: 100 g of jam to be produced

Therefore, Amount of Jam, m3=100 g

Solid二x3 Rasidue: Watu 3 85 uott. wati Solid 1-/3-3 vohee m is the Mass o Sthauoberuiss ma is the Mass m3 is th Amount o Jam

Assumption: Steady state process with no reaction

Given: The crushed strawberries and sugar is mixed in a 4/5 mass ratio.

\frac{m_{1}}{m_{2}}=\frac{4}{5}

{m_{2}}=\frac{5}{4}{m_{1}}

m2= 1.25 m1 - Equation 1

Overall Mass Balance is Mass In=Mass Out i.e., m1+m2=100+m4 - Equation 2

A material balance for the solid is 0.15m1=x3*m3 (or) 0.15m1=100g*x3 - Equation 3

A material balance for water is 0.85m1=1/3 m3+m4

0.85m1= (1/3*100g)+m4

m4=0.85m1-(1/3*100g) - Equation 4

We have Equation 2 as m1+m2=100+m4

Substitute for m2 from Equation 1,

m1+1.25m1=100+m4

Substitute for m4 from Equation 4,

m1+1.25m1=100+(0.85m1-(1/3*100g))

Solving for m1, we get

2.25m1= 100+0.85m1-33.33

1.4m1= 66.67

m1= 47.62 g

To find m2: Substitute the value of m1 in Equation 1

m2= 1.25 m1= 1.25*47.62g

m2= 59.52 g

To find m4: Substitute the value of m1 in Equation 4,

m4=0.85m1-(1/3*100g)

m4= (0.85*47.62g)-(1/3*100g)

m4= 7.14 g

To find x3: Substitute the value of m1 in Equation 3

0.15m1=100g*x3

0.15*47.62g=100g*x3

x3= 0.0714

Amount of Strawberries required is m1=47.62 g

Amount of Water Evaporated is m4= 7.14 g

2.i) Quencher: A quencher is typically a spray system used to quench hot gases i.e., to lower the temperature of gases when it is higher than the ambient temperature. As the gases can evaporate large amounts of scrubbing liquid, quenching is done in order to retain the liquid. This avoids damage to the internal parts of scrubber. Quencher is designed using the principle that the operational efficiency increases with increase in liquid-gas contact.  

ii) Decanter: Decantation is a process of separating liquid and solids from a suspension or separation of immisicible liquids. The vessel used for this separation is called the decanter. In decanter, separation is based on the densities or gravity of the solids and liquids.

iii) Turbine: A flowing fluid has potential and kinetic energy. Turbine is the rotary device used to extract work from flowing fluids and covert it into useful work. Steam Turbine is used in thermal plants as it extracts thermal energy from the steam. Water Turbine is used to extract hydro energy from flowing water.

iv) Humidifier: Humidification is the process of increasing the moisture in a given space. The device used for this is called the humidifier. Humidifier is used to maintain specific level of humidity for a process to preserve the properties of the material. It is also used to prevent accumulation of static electricity.

v) Partial Condenser: A partial condenser is used in separation process (like distillation). It condenses the vapor partially and this condensed liquid is sent back to the distillation column. It is used when total condensation is not possible due to lower boiling point of distillate.